In the two-sector economy, Y = Yd. Prove Y = Yd.


1
Expert's answer
2021-05-25T11:08:59-0400

According to Keynes, a two sector economy is a closed economy comprising of household and firms only, as sector players. there no government influence and hence taxation on the income for both players.

"\therefore" Y = Yd where Y is the GDP = C + I and Yd = Y - T or C + I -T

C is Consumers expenditurE, I is Investment expenditure by firms and T is taxes levied by the government, which in this case of two sector economy is zero (0)

Hence Y = Y-T or C + I = C + I -0

"\because" At equilibrium, C + I = C + I